Please help me how do you solve this problem?
1/2a+4b=11
a-5b=-4
I have to do this problem by using elimination

Answer by kapilsinghi(68)

1/2a+4b=11 --------1
a-5b=-4 --------2
multiplying 1 by 2
a+8b=22 --------3
a-5b=-4 --------4
subtracting 4 from 3
13b=26
b = 2
subtituting value of b in 2
a - 10 = -4
a = 6

kapilsinghi123@gmail.com